A chef is using a mixture of two brands of Italian dressing. The first brand contains 5% vinegar, and the second brand contains 15% vinegar. The chef wants to make 280 milliliters of a dressing that is 9% vinegar. How much of each brand should she use?
Answer by TimothyLamb(4379)

---
5x + 15y = 9*280
x + y = 280
---
5x + 15y = 2520
x + y = 280
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 168 ml of 5%
y= 112 ml of 15%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php